# Efficient Diffusion Model for Image Restoration by Residual Shifting

### v1

@@ Abstract
While diffusion-based image restoration (IR) methods have achieved remarkable success, they are still limited by the low inference speed attributed to the necessity of executing hundreds or even thousands of sampling steps. Existing acceleration sampling techniques, though seeking to expedite the process, inevitably sacrifice performance to some extent, resulting in over-blurry restored outcomes. To address this issue, this study proposes a novel and efficient diffusion model for IR that significantly reduces the required number of diffusion steps. Our method avoids the need for post-acceleration during inference, thereby avoiding the associated performance deterioration. Specifically, our proposed method establishes a Markov chain that facilitates the transitions between the high-quality and low-quality images by shifting their residuals, substantially improving the transition efficiency. A carefully formulated noise schedule is devised to flexibly control the shifting speed and the noise strength during the diffusion process. Extensive experimental evaluations demonstrate that the proposed method achieves superior or comparable performance to current state-of-the-art methods on three classical IR tasks, namely image super-resolution, image inpainting, and blind face restoration, even only with four sampling steps. Our code and model are publicly available at https://github.com/zsyOAOA/ResShift.
@@ I Introduction
Image restoration (IR) is a fundamental problem in low-level vision, aiming at recovering a high-quality (HQ) image given its corresponding low-quality (LQ) counterpart. It encompasses various sub-tasks upon the type of degradation, including but not limited to image denoising [1, 2], image super-resolution [3, 4] and image inpainting [5, 6]. Notably, the degradation model is usually complicated in some cases, e.g., real-world super-resolution, making the IR problem severely ill-posed and challenging.
@@
The diffusion model [7], a newly emerged generative model, has demonstrated remarkable success in the realm of image generation [8, 9], surpassing the previous GAN-based approaches [10, 11]. The main idea of diffusion models is to construct a hidden Markov chain that gradually degrades an image to white Gaussian noise in the forward process, and then approximate the reverse process using a deep neural network to generate this image. Attributed to its powerful generative capability, diffusion model has also exhibited great promise in solving the IR tasks, including image denoising [12, 13], deblurring [14, 15], inpainting [16, 17, 18, 19], colorization [20, 21, 22], among others. The exploration of the potential of diffusion models in IR still remains an active and ongoing research pursuit.
@@
Existing diffusion-based IR methods can be broadly categorized into two classes. One straightforward approach [23, 24, 21, 25, 26, 27] is to insert the LQ image into the input of current diffusion model, e.g., DDPM [8], as a condition. Then, the model is retrained specifically for the task of IR. After training, this model is able to generate a desirable HQ image from Gaussian noise and the observed LQ image via the reverse sampling process. Another popular way, as explored in [28, 29, 30, 31, 32, 33, 34], attempts to harness a pre-trained unconditional diffusion model as a prior to facilitate the IR problem. The primary idea involves modifying the reverse sampling procedure to align the generated result with the given LQ observation via the degradation model at each step. Unfortunately, both strategies inherit the Markov chain underlying DDPM, which can be inefficient during inference, i.e., often necessitating hundreds or even thousands of sampling steps. Although some acceleration techniques [35, 36, 37] have been developed to reduce the sampling steps in inference, they inevitably lead to a significant drop in performance, resulting in over-smooth results as shown in Fig. 1(e)-(j), in which the DDIM [36] algorithm is employed to speed up the inference. Therefore, there is a need to devise a new diffusion model tailored for IR, capable of achieving a harmonious balance between efficiency and performance, without sacrificing one for the other.
@@
Let us revisit the diffusion model within the context of image generation. In the forward process, it constructs a Markov chain to gradually transform the observed data into a pre-specified prior distribution, typically a standard Gaussian distribution, over a large number of steps. Conversely, the reverse process trains a deep neural network to approximate the reverse trajectory of this Markov chain. The well-trained neural network facilitates the random image generation by drawing samples from the reverse Markov chain, initiating at the Gaussian distribution. While the Gaussian distribution is well-suited for the task of image generation, its optimality is questioned in the domain of IR, where the LQ images are available. In this paper, we argue that a reasonable diffusion model for IR should start from a prior distribution centered around the LQ image, enabling an iterative recovery of the HQ image from its LQ counterpart instead of Gaussian white noise. Notably, this design not only aligns with the specifics of IR but also holds the potential to reduce the requisite number of diffusion steps for sampling, thereby enhancing overall inference efficiency.
@@
Figure 1: Qualitative comparisons on one typical real-world example of the proposed method and recent state of the arts, including RealESRGAN [38], BSRGAN [39], SwinIR [40], LDM [25], and StableSR [29]. As for the diffusion-based approaches, namely LDM, StableSR, and our proposed method, we annotate the number of sampling steps with the format of “Method-A” for more intuitive visualization, where “A” denotes the number of sampling steps. Note that LDM and StableSR both contain 1000 diffusion steps in training and are accelerated to “A” steps using DDIM [36] during inference. Please zoom in for a better view.
@@
Following the aforementioned motivation, we propose an efficient diffusion model involving a shorter Markov chain for the transition between the HQ image and its corresponding LQ one. The initial state of the Markov chain converges towards an approximate distribution of the HQ image, while the final state converges towards an approximate distribution of the LQ image. To achieve this, we carefully design a transition kernel that gradually shifts the residual information between the HQ/LQ image pair. This approach exhibits superior efficiency in comparison to existing diffusion-based IR methods, given its ability to expeditiously transfer the residual information in several steps. Moreover, our design also allows for an analytical and concise expression for the evidence lower bound, thereby simplifying the formulation of the optimization objective for training purposes. Building upon the constructed diffusion kernel, we further develop a highly flexible noise schedule that controls the speed of residual shifting and the strength of the added noise in each step. This schedule facilitates a fidelity-realism trade-off of the recovered results by tuning its hyper-parameters.
@@
  • The proposed method is a general diffusion-based framework for IR, being able to handle various IR tasks. This study has thoroughly substantiated its effectiveness and superiority on three typical and challenging IR tasks, namely image super-resolution, image inpainting, and blind face restoration.
In summary, this work endeavors to formulate an efficient diffusion model customized for IR, with the intention of breaking down the limitation of prevailing approaches on inference efficiency. A preliminary version of this work has been published in NeurIPS 2023 [41], focusing only on the task of image super-resolution. This study makes substantial improvements in both model design and empirical evaluation across diverse IR tasks compared with the conference version. Concretely, we incorporate the perceptual loss into the model optimization and substitute the self-attention layer with Swin transformer in the network architecture. The former modification can further reduce the diffusion steps from 15 to 4, and the latter endows our model with graceful adaptability to handle arbitrary resolutions during inference.
The remainder of the manuscript is organized as follows: Sec. II introduces the related work. Sec. III presents our designed diffusion model for IR. In Sec. IV and Sec. V, extensive experiments are conducted to evaluate the performance of our proposed method on the task of image super-resolution and image inpainting, respectively. Sec. VI finally concludes the paper.
@@ II-B Diffusion-based Image Restoration Approaches
Inspired by principles from non-equilibrium statistical physics, Sohl-Dickstein et al. [7] proposed the diffusion model to fit complex distributions. Subsequent advancements by Ho et al. [8] and Song et al. [20] further improve its theoretical foundation by integrating denoising score matching and stochastic differential equation, thereby achieving impressive success in image generation [25]. Owing to its powerful generative capability, diffusion models have also found successful applications in the field of IR. Next, we provide a comprehensive overview of recent developments in diffusion-based IR methods.
The most straightforward solution to solve the IR problem using diffusion models is to introduce the LQ image as an addition condition in each timestep. Pioneering this approach, Saharia et al. [23] have successfully trained a diffusion model for image super-resolution. Subsequent studies [21, 14, 27] further expanded upon this approach, exploring its applicability in image deblurring, colorization, and denoising. To circumvent the resource-intensive process of training from scratch, an alternative strategy involves harnessing a pre-trained diffusion model to facilitate IR tasks. Numerous investigations, such as [18, 24, 28, 17, 32, 33, 15], reformulated the reverse sampling procedure of a pre-trained diffusion model into an optimization problem by incorporating the degradation model, enabling solving the IR problem in a zero-shot manner. Most of these methods, however, cannot deal with the blind IR problem, as they rely on a pre-defined degradation model. In contrast, some other works [29, 75, 30] directly introduced a trainable module that takes the LQ image as input. This module modulates the feature maps of the pre-trained diffusion model, steering it toward the direction of generating a desirable HQ image. Such a paradigm eliminates the reliance on a degradation model in the test phase, rendering it capable of handling the blind IR tasks.
@@ III-A Model Design
The optimization for $\bm{\theta}$ is achieved by minimizing the negative evidence lower bound, namely,
$\min_{\bm{\theta}}\sum_{t}D_{\text{KL}}\left[q(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0},\bm{y}_{0})\|p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t},\bm{y}_{0})\right],$ (6)
@@
$\min_{\bm{\theta}}\sum\nolimits_{t}w_{t}\|f_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)-\bm{x}_{0}\|_{2}^{2},$ (10)
where $w_{t}=\frac{\alpha_{t}}{2\kappa^{2}\eta_{t}\eta_{t-1}}$. In practice, we empirically find that the omission of weight $w_{t}$ results in a notable performance improvement, aligning with the conclusion in [8].
Perceptual Regularization. As presented above, our proposed method facilitates an iterative restoration process starting from the LQ image, in contrast to prior methods that initialize from Gaussian noise. This approach effectively reduces the number of diffusion steps. A comprehensive experimental analysis in Sec. IV-B substantiates that our method yields promising results with a mere 15 sampling steps, demonstrating a notable acceleration compared to established methodologies [25, 29].
Unfortunately, attempts at further acceleration, particularly with fewer than 5 sampling steps, tend to produce over-smooth results. This phenomenon is primarily attributed to that the $L_{2}$-based loss in Eq. (10) favors the prediction of an average over plausible solutions. To overcome this limitation, we introduce the perceptual regularization [77] to impose an additional constraint on the solution space, namely,
$\min_{\bm{\theta}}\sum\nolimits_{t}\|f_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)-\bm{x}_{0}\|_{2}^{2}+\lambda l_{p}\left(f_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t),\bm{x}_{0}\right),$ (11)
where $l_{p}(\cdot,\cdot)$ denotes the pre-trained LPIPS metric, $\lambda$ is a hyper-parameter controlling the relative importance of these two constraints. This simple solution effectively curtails the sampling trajectory to fewer steps, e.g., 4 steps in this study, as shown in Sec. IV-B, while concurrently maintaining superior performance.
@@ III-B Noise Schedule
The proposed noise schedule exhibits a high degree of flexibility in three key aspects. First, in the case of small values of $\kappa$, the final state $\bm{x}_{T}$ converges towards a perturbation around the LR image, as illustrated in Fig. 3(c)-(d). Compared to the diffusion process ended at Gaussian noise, this design significantly shortens the length of the Markov chain, thereby improving the inference efficiency. Second, the hyper-parameter $p$ provides precise control over the shifting speed, enabling a fidelity-realism trade-off in the SR results as analyzed in Sec. IV-B. Third, by setting $\kappa=40$ and $p=0.8$, our method achieves a diffusion process that remarkably resembles to LDM [25]. This is clearly demonstrated by the visual results of the diffusion process presented in Fig. 3(e)-(f), and further supported by the comparisons on the relative noise strength as shown in Fig. 3(g).
@@
Figure 4: Visual comparison of two different models containing some self-attention layers (denoted as model-1) or Swin transformers (denoted as model-2). (a1) and (a2): zoomed LQ images with resolutions of $64\times 64$ or $128\times 128$. (b1) and (b2): super-resolved results by model-1. (c1) and (c2): visualized attention maps extracted from the first self-attention layer of model-1. Note that these visualized results are obtained by first calculating the first principal component of the attention map via PCA and then reshaping it to the targeted size. In the left-upper corner, we annotate the average entropy values of these attention maps. (d1) and (d2): super-resolved results by model-2.
@articleliang2024control, title=Control Color: Multimodal Diffusion-based Interactive Image Colorization, author=Liang, Zhexin and Li, Zhaochen and Zhou, Shangchen and Li, Chongyi and Loy, Chen Change, journal=arXiv preprint arXiv:2402.10855, year=2024
III-C Discussion on Arbitrary Resolution

### v2

@@ Abstract
While diffusion-based image restoration (IR) methods have achieved remarkable success, they are still limited by the low inference speed attributed to the necessity of executing hundreds or even thousands of sampling steps. Existing acceleration sampling techniques, though seeking to expedite the process, inevitably sacrifice performance to some extent, resulting in over-blurry restored outcomes. To address this issue, this study proposes a novel and efficient diffusion model for IR that significantly reduces the required number of diffusion steps. Our method avoids the need for post-acceleration during inference, thereby avoiding the associated performance deterioration. Specifically, our proposed method establishes a Markov chain that facilitates the transitions between the high-quality and low-quality images by shifting their residuals, substantially improving the transition efficiency. A carefully formulated noise schedule is devised to flexibly control the shifting speed and the noise strength during the diffusion process. Extensive experimental evaluations demonstrate that the proposed method achieves superior or comparable performance to current state-of-the-art methods on four classical IR tasks, namely image super-resolution, image inpainting, blind face restoration, and image deblurring, even only with four sampling steps. Our code and model are publicly available at https://github.com/zsyOAOA/ResShift.
@@ I Introduction
Image restoration (IR) is a critical challenge in the field of low-level vision, with the goal of recovering a high-quality (HQ) image from its corresponding low-quality (LQ) variant. This challenge can be further divided into different sub-tasks upon its degradation model, including image super-resolution [1, 2], image deblurring [3, 4], and image inpainting [5, 6], among others [7, 8]. Particularly, the degradation models encountered in practical scenarios, such as those in real-world super-resolution, often exhibit significant complexity, rendering the IR problem severely ill-posed and challenging to address.
@@
The diffusion model [9] has revolutionized the traditional paradigm of image generation based on Generative Adversarial Networks (GANs) [10, 11], further advancing the field of image synthesis [12, 13]. This approach leverages a hidden Markov chain to progressively corrupt an image into white Gaussian noise through a forward diffusion process, and subsequently employs a deep neural network to approximate the reverse process for image reconstruction. Attributed to its powerful generative capability, the diffusion model has shown considerable potential in addressing various IR tasks, including image denoising [14, 15], deblurring [4, 16], inpainting [17, 18, 19, 20], colorization [21, 22, 23]. The exploration of diffusion models’ capabilities in IR still remains an active and promising area of research.
@@
In this study, we categorize recent diffusion-based IR methods into two main approaches. The first approach [24, 22, 25, 26, 27] directly incorporates the LQ image into the input of a current diffusion model, such as DDPM [12], as a condition, and then retrain this model specifically for the IR task. Once trained, this model can generate the desirable HQ image from Gaussian noise and the observed LQ image through the reverse sampling process. The second approach, as explored in [28, 29, 30, 31, 32, 33, 34, 35], adopts a pre-trained unconditional diffusion model as a prior to address the IR problem. This method modifies the reverse sampling procedure to align the generated outputs with the given LQ observations by incorporating the degradation model at each iteration. However, both strategies are limited by the inherent Markov chain structure of DDPM, which often leads to inefficiencies during inference, requiring hundreds or even thousands of sampling steps. While recent advancements [36, 37, 38] have introduced acceleration techniques to reduce sampling steps, these methods inevitably result in a significant performance drop, as evidenced by the overly smooth results shown in Fig. 1(i)-(k). Thus, there is a need to devise a new diffusion model specifically designed for IR that achieves an optimal trade-off between efficiency and performance, without sacrificing one for the other.
@@
In the domain of image generation, diffusion models progressively convert the observed data into a pre-determined prior distribution, often a standard Gaussian distribution, through a Markov chain over numerous steps. The forward diffusion process constructs this Markov chain, while the reverse process involves training a deep neural network to approximate the inverse trajectory of the Markov chain. The trained neural network can generate images randomly by sampling from the reverse Markov chain, initiating at the Gaussian distribution. Although the Gaussian distribution is well-suited for image generation, its optimality is questioned for IR, where LQ images are available as extra information. In this paper, we argue that a reasonable diffusion model for IR should start from a prior distribution centered around the LQ image, enabling an iterative recovery of the HQ image from its LQ counterpart instead of Gaussian white noise. This approach not only aligns more closely with the characteristics of IR but also holds the potential to reduce the number of diffusion steps for sampling, thereby improving inference efficiency.
@@
Figure 1: Qualitative comparisons on one typical real-world example of the proposed method and recent state-of-the-arts, including RealESRGAN [39], BSRGAN [40], SwinIR [41], LDM [25], StableSR [30], and CCSR [42]. As for the diffusion-based approaches and our proposed method, we annotate the number of sampling steps with the format of “Method-A” for more intuitive visualization, where “A” denotes the number of sampling steps.
@@
In light of the preceding motivation, we introduce an efficient diffusion model characterized by a shorter Markov chain transferring between the HQ and LQ images. The Markov chain’s initial state converges towards an approximate distribution of the HQ image, while the final state approximates the LQ image distribution. This is achieved through the design of a transition kernel that incrementally shifts residual information between the HQ and LQ image pair. Our method exhibits superior efficiency beyond existing diffusion-based IR methods, due to its capacity to rapidly transfer residual information across a limited number of steps. Moreover, this design also allows for an analytical and concise expression of the evidence lower bound (ELBO), thereby simplifying the formulation of the optimization objective for training. Beyond the traditional ELBO, we empirically find that introducing a perceptual regularizer can further reduce the diffusion steps during training, and thus improve the inference efficiency. Building upon the constructed diffusion kernel, we develop a highly flexible noise schedule that controls the rate of residual transfer and the intensity of the added noise at each step. This schedule provides a mechanism for balancing the fidelity and realism of the recovered images by tuning its hyper-parameters.
@@
  • Based on the traditional diffusion Unet, we propose to substitute its self-attention layers with Swin Transformer blocks to enhance its capability in handling images with varying resolutions.
  • The proposed method is a general diffusion-based framework for IR and capable of handling various IR tasks. This study has thoroughly substantiated its effectiveness and superiority on four typical and challenging IR tasks, namely image super-resolution, image inpainting, blind face restoration, and image deblurring.
In summary, our work formulates an efficient diffusion model tailored for IR, overcoming the limitation of prevailing approaches on inference efficiency. A preliminary version of this work has been published in NeurIPS 2023 [43], focusing only on the task of image super-resolution. This study makes substantial improvements in both model design and empirical evaluation across diverse IR tasks compared with the conference version. Concretely, we incorporate the perceptual loss into the model optimization and substitute the self-attention layer with shifted window-based self-attention presented in Swin Transformer [44] in the network architecture. The former modification can further reduce the diffusion steps from 15 to 4, and the latter endows our model with graceful adaptability to handle arbitrary resolutions during inference.
The remainder of the manuscript is organized as follows: Section II introduces the related work. Section III presents our designed diffusion model for IR. In Section IV and Section V, extensive experiments are conducted to evaluate the performance of our proposed method on the task of image super-resolution and image inpainting, respectively. Section VI finally concludes the paper.
@@ II-B Diffusion-based Image Restoration Approaches
Inspired by principles from non-equilibrium statistical physics, Sohl-Dickstein et al. [9] proposed the diffusion model to fit complex distributions. Subsequent advancements by Ho et al. [12] and Song et al. [21] further improve its theoretical foundation by integrating denoising score matching and stochastic differential equation, thereby achieving impressive success in image generation [25, 78]. Owing to its powerful generative capability, diffusion models have also found successful applications in the field of IR. Next, we provide a comprehensive overview of recent developments in diffusion-based IR methods.
The most straightforward solution to solve the IR problem using diffusion models is to introduce the LQ image as an addition condition in each timestep. Pioneering this approach, Saharia et al. [24] have successfully trained a diffusion model for image super-resolution. Subsequent studies [22, 4, 27] further expanded upon this approach, exploring its applicability in image deblurring, colorization, and denoising. To circumvent the resource-intensive process of training from scratch, an alternative strategy involves harnessing a pre-trained diffusion model to facilitate IR tasks. Numerous investigations, such as [19, 29, 28, 18, 33, 34, 16, 79], reformulated the reverse sampling procedure of a pre-trained diffusion model into an optimization problem by incorporating the degradation model, enabling solving the IR problem in a zero-shot manner. Most of these methods, however, cannot deal with the blind IR problem, as they rely on a pre-defined degradation model. In contrast, some other works [30, 80, 31, 81, 82] directly introduced a trainable module that takes the LQ image as input. This module modulates the feature maps of the pre-trained diffusion model, steering it toward the direction of generating a desirable HQ image. Such a paradigm eliminates the reliance on a degradation model in the test phase, rendering it capable of handling the blind IR tasks.
@@ III-A Model Design
Algorithm 1 Training
0:  Degradation model $\mathcal{D}(\cdot)$, high-quality dataset $\mathcal{T}$
1:  repeat
2:     $\bm{x}_{0}\sim\mathcal{T}$, $\bm{y}_{0}=\mathcal{D}(\bm{x}_{0})$
3:     $t\sim\text{Uniform}\left(\{1,\cdots,T\}\right)$
4:     $\bm{x}_{t}\sim q(\bm{x}_{t}|\bm{x}_{0},\bm{y}_{0})$
5:     Take gradient descent step on $\nabla\mathcal{L}_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)$
6:  until converged
@@
The optimization for $\bm{\theta}$ is achieved by minimizing the following negative ELBO, i.e.,
$\sum_{t}D_{\text{KL}}\left[q(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0},\bm{y}_{0})\|p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t},\bm{y}_{0})\right],$ (6)
@@
Algorithm 2 Sampling
0:  Low-quality image $\bm{y}$
1:  $\bm{x}_{T}\sim\mathcal{N}(\bm{x}_{T};\bm{y},\kappa^{2}\eta_{T}\bm{I})$
2:  for $t=T,\cdots,1$ do
3:     $\bm{\epsilon}\sim\mathcal{N}(\bm{\epsilon};\bm{0},\bm{I})$ if $t>1$ else $\bm{\epsilon}=\bm{0}$
4:     $\bm{\mu}=\frac{\eta_{t-1}}{\eta_{t}}\bm{x}_{t}+\frac{\alpha_{t}}{\eta_{t}}f_{\bm{\theta}}(\bm{x}_{t},\bm{y},t)$
5:     $\bm{x}_{t-1}=\bm{\mu}+\kappa\sqrt{\frac{\eta_{t-1}\alpha_{t}}{\eta_{t}}}\bm{\epsilon}$
6:  end for
7:  return  $\bm{x}_{0}$
@@
$\mathcal{L}_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)=\sum\nolimits_{t}w_{t}\|\hat{\bm{x}}_{0}^{t}-\bm{x}_{0}\|_{2}^{2},$ (10)
where $w_{t}=\frac{\alpha_{t}}{2\kappa^{2}\eta_{t}\eta_{t-1}}$, $\hat{\bm{x}}_{0}^{t}=f_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)$. In practice, we empirically find that the omission of weight $w_{t}$ results in a notable performance improvement, aligning with the conclusion in [12]. And the detailed training process is summarized in Algorithm 1. After training, we can generate the desirable HQ result following Algorithm 2 for any LQ testing image.
Perceptual Regularization. As presented above, our proposed method facilitates an iterative restoration process starting from the LQ image, in contrast to prior methods that initialize from Gaussian noise. This approach effectively reduces the number of diffusion steps. The comprehensive experimental analysis in our conference paper [43] has substantiated that the proposed method yields promising results with a mere 15 sampling steps, demonstrating a notable acceleration compared to established methodologies [25, 30].
Unfortunately, attempts at further acceleration, particularly with fewer than 5 sampling steps, tend to produce over-smooth results. This phenomenon is primarily attributed to that the $L_{2}$-based loss in Eq. (10) favors the prediction of an average over plausible solutions. To overcome this limitation, we introduce an additional perceptual regularization [84] on Eq. (10) to further constrain the solution space, namely,
$\mathcal{L}_{\bm{\theta}}(\bm{x}_{t},\bm{y}_{0},t)=\sum\nolimits_{t}\|\hat{\bm{x}}_{0}^{t}-\bm{x}_{0}\|_{2}^{2}+\lambda l_{p}\left(\hat{\bm{x}}_{0}^{t},\bm{x}_{0}\right),$ (11)
where $l_{p}(\cdot,\cdot)$ denotes the pre-trained LPIPS metric, $\lambda$ is a hyper-parameter controlling the relative importance of these two constraints. This solution effectively curtails the sampling trajectory to fewer steps, e.g., 4 steps in this study, while concurrently maintaining superior performance.
@@ III-B Noise Schedule
The proposed noise schedule exhibits a high degree of flexibility in three key aspects. First, in the case of small values of $\kappa$, the final state $\bm{x}_{T}$ converges towards a perturbation around the LQ image, as illustrated in Fig. 3(c)-(d). Compared to the diffusion process ended at Gaussian noise, this design significantly shortens the length of the Markov chain, thereby improving the inference efficiency. Second, the hyper-parameter $p$ provides precise control over the shifting speed, enabling a fidelity-realism trade-off in the SR results as analyzed in Sec. IV-B. Third, by setting $\kappa=40$ and $p=0.8$, our method achieves a diffusion process that degenerates into LDM [25]. This is clearly demonstrated by the visual results of the diffusion process presented in Fig. 3(e)-(f), and further supported by the comparisons on the relative noise strength as shown in Fig. 3(g).
@@
Figure 4: Visual comparison of two different models containing some self-attention layers (denoted as model-1) or Swin Transformers (denoted as model-2). (a1) and (a2): zoomed LQ images with resolutions of $64\times 64$ or $128\times 128$. (b1) and (b2): super-resolved results by model-1. (c1) and (c2): visualized attention maps extracted from the first self-attention layer of model-1. Note that these visualized results are obtained by first calculating the first principal component of PCA of the attention map and then reshaping it to the targeted size. In the left-upper corner, we annotate the entropy value of these attention maps. (d1) and (d2): super-resolved results by model-2.
III-C Relation to Flow Matching
Flow matching [86], also known as Rectified flow [87], is another advanced framework beyond diffusion models, focusing on finding the optimal transport map from one distribution to another. In this section, we present an alternative formulation of our proposed method through flow matching, offering a novel perspective on its theoretical foundation.
We first introduce two important definitions, namely the probability density path $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$, which is a time-dependent probability density function satisfying $\int p_{t}(x)\operatorname{d\!}x=1$, and a time-dependent vector field $v_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. Given the data points $\bm{x}\in\mathbb{R}^{d}$, the vector field $v_{t}$ can be used to construct a time-dependent diffeomorphic map, called a flow $\phi_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, which is defined by the following ordinary differential equation:
${\color[rgb]{0,0,0}\frac{\operatorname{d\!}}{\operatorname{d\!}t}\phi_{t}(\bm{x})}$ ${\color[rgb]{0,0,0}=v_{t}(\phi_{t}(\bm{x})),}$ (14a)
${\color[rgb]{0,0,0}\phi_{0}(\bm{x})}$ ${\color[rgb]{0,0,0}=\bm{x}.}$ (14b)
Chen et al. [88] proposed to parameterize the vector field $v_{t}$ as a deep neural network $v_{t}(\cdot;\theta)$ with parameter $\theta$, leading to a deep parametric model of the flow $\phi_{t}$, called continuous normalizing flow (NF). In image generation, NF is often used to model the transport map between one simple known distribution $q_{0}$, typically Gaussian, and the data distribution $q_{1}$. Lipman et al. [86] developed the conditional flow matching technique that defines conditional probability path as follows:
$p_{t}(\bm{x}|\bm{x}_{1})=\mathcal{N}(\bm{x}|\mu_{t}(\bm{x}_{1}),\sigma_{t}(\bm{x}_{1})^{2}\bm{I}),$ (15)
where $\bm{x}_{1}\sim q_{1}(\bm{x})$. Furthermore, the corresponding flow $\phi_{t}$ is specified in a simple format:
$\phi_{t}(\bm{x})=\sigma_{t}(\bm{x}_{1})\bm{x}+\mu_{t}(\bm{x}_{1}),$ (16)
where $\bm{x}\sim q_{0}(x)$. This work provides a general framework with a close relationship to diffusion models and optimal transport theory, and more details can be found in [86, 89].
Even though our proposed method is formulated upon the diffusion model, it corresponds to a conditional flow between the LQ image distribution $q_{0}$ and the HQ image distribution $q_{1}$, specifically designed for IR. For a given image pair $\bm{y}_{0}$ and $\bm{x}_{0}$ from $q_{0}$ and $q_{1}$ respectively, the underlying probability path $p_{t}$ of our method can be expressed as:
$p_{t}(\bm{x}|\bm{x}_{0})=\mathcal{N}(\bm{x}|\mu_{t}(\bm{x}_{0}),\sigma_{t}(\bm{x}_{0})^{2}\bm{I}),$ (17)
where
$\mu_{t}(\bm{x}_{0})=\eta_{t}\bm{x}_{0}+(1-\eta_{t})\bm{y}_{0},~{}\sigma_{t}(\bm{x}_{0})=\kappa\sqrt{1-\eta_{t}},$ (18)
and $\eta_{t}$ is defined in Eq. (12). The conditional flow $\phi_{t}$ is a linear interpolation between the LQ and HQ images followed with a noise disturbance, i.e.,
$\phi(\bm{y}_{0})=\eta_{t}\bm{x}_{0}+(1-\eta_{t})\bm{y}_{0}+\kappa\sqrt{1-\eta_{t}}\bm{\xi},$ (19)
where $\bm{y}_{0}\sim q_{0}(\bm{y}_{0})$, $\bm{\xi}\sim\mathcal{N}(\bm{0},\bm{I})$. This intuitive and straightforward path provides a rapid transport map between LQ and HQ distributions, thereby improving the sampling inference significantly, aligning with the conclusions in [86, 87].
III-D Discussion on Arbitrary Resolution